two identical objects are placed 0.2 m apart and experience gravitational force of 2.0 *10-9 N. find the masses of spheres.
1
Expert's answer
2021-03-23T11:08:15-0400

If the size of the objects is small compared to the distance between them, then


"F=G\frac{m\cdot m }{r^2}\to m=\sqrt{\frac{Fr^2}{G}}=\sqrt{\frac{2\cdot10^{-9}\cdot 0.2^2}{6.674\cdot10^{-11}}}\approx1.1\ kg" . Answer